OpenStudy (anonymous):

Write an explicit rule for t(n) of the geometric sequence below: 2, 3/2, 9/8, 27/32 .....

OpenStudy (anonymous):

2* (3/4)=3/2

OpenStudy (anonymous):

so thats the rule

OpenStudy (anonymous):

no

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

\[a_{n}=ar^{n-1}\]is the general formula for geometric sequences. Here the "common ratio" is \[r=\frac{\frac{3}{2}}{2}=\frac{3}{4}\]The first term "a" is 2. So we have:\[t(n)=2(\frac{3}{4})^{n-1}\]where\[n=1,2,3...\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so will the rule be n(t)=1,2,3

OpenStudy (anonymous):

no. The rule is,\[t(n)=2(\frac{3}{4})^{n-1}\]where,\[n=1,2,3...\]

OpenStudy (anonymous):

so the rule is 2(3/4)n−1

OpenStudy (anonymous):

(n-1) is the exponent on the (3/4) term. It is:\[(\frac{3}{4})^{n-1}\]NOT,\[\frac{3}{4}(n-1)\]

OpenStudy (anonymous):

so it is (3/4)^n-1

OpenStudy (anonymous):

For the (3/4) term, yes. look at my previous post for the whole equation.

OpenStudy (anonymous):

so 2(3/4)^n-1

OpenStudy (anonymous):

2(3/4)^{n-1}

OpenStudy (anonymous):

Notice:\[2(\frac{3}{4})^{n-1}\neq 2(\frac{3}{4})^n-1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2(3/4)^{n-1} it is like this

OpenStudy (anonymous):

yep. So the first term, n=1, gives, 2(3/4)^{1-1}=2(3/4)^(0)=2, as expected. And the second term in the sequence is...n=2...... 2(3/4)^{2-1}=3/2....etc...

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

cool